Solve:
3w-2x+4y-z=-2
5w+3x-2y+7z=-36
2w-5x+6y-2z=2
-4w+6x-8y+9z=-26

Answer by Alan3354(69443)

3w-2x+4y-z=-2
5w+3x-2y+7z=-36
2w-5x+6y-2z=2
-4w+6x-8y+9z=-26
---------------
I have an Excel sheet that does these using determinants.
w = -2
x = 1
y = 1/2
z = -4
------
If you want info on how to do them, email me via the thank you note.